OpenStudy (anonymous):

sin theta= 24/25 : find cos (2theta)

OpenStudy (anonymous):

you need to find the cosine of theta first do you know how to do that?

OpenStudy (anonymous):

@satellite73 0<theta<pi/2

OpenStudy (anonymous):

Alternatively, find theta = arc sin (24/25), then get cos (2 theta). Assuming numerical answer is sought.

OpenStudy (anonymous):

cos 2x = 1 - 2sin^2 x sin x = 24/25 , -> sin^2 x = (24/25)^2 = (576/625) cos 2x = 1 - 2(576/625) = -527/625 =-0.84